# A Diagnostic to Find and Help Combat Positivity Issues – with a Focus on Continuous Treatments

### v1

@@ 5.2 Single Time Point Analysis
At first, in order to simplify the problem at hand, the question is reduced to a single time point target. Therefore the variable set reduces to the initial weight, general adherence, and, if effect modification is of interest, the metabolization group. In addition, the mid-dose concentration at weeks 6 and 36 are the treatment of interest, which we assume to be the same at both time points for simplicity. The target is the first viral failure measured at week 36. The estimand is therefore $P(VF_{t=36}^{EVF_{t=6}=EVF_{t=36}=a}=1)$ for $a\in[0,6]$, so the dose-response curve of viral failure in week 36 had the mid-dose concentration of efavirenz been $a$ in both weeks 6 and 36. In the following the diagnostic is used in order to analyze how much positivity/sparsity issues influence the estimation of a CDRC. This is done first for a general causal estimate and then in a second analysis for a causal estimate which includes the estimation of effect modification by the metabolization group.
@@ 5.2.1 Causal Analysis
Figure 5: The kernel density with its chosen hyperparameters for an $o^{int}$, with weight 3 and mid-dose concentration 5
@@
Before the diagnostic can be used the hyperparameters have to be fixed. We chose a Gaussian kernel as we did in the simulation study above. Next, the hyperparameters controlling how wide or narrow the kernel is in every direction need to be set. Specifically, the question “how far do we think we can extra- and interpolate in each dimension?” has to be asked. For the binary variable adherence, we assume that no extrapolation is possible. For the other adjustment variable weight, we assume that an observation which shares every other variable, but deviates in weight by one standard deviation should get a diagnostic value of 0.5, meaning it would be treated as half an observation of support. As mid-dose concentration is the treatment variable, we set a more narrow shape in that direction with only half a standard deviation leading to a half observation of support. We can illustrate how that kernel will influence the diagnostic by assuming we are interested in ascribing a concentration of 5 to a person with weight score 3, see Figure 5. If we only consider these two dimensions, this leads to a diagnostic value of roughly 4.3 which is mostly due to the 5 points close by. Plots like this can also help researchers gain a deeper understanding on what certain hyperparameter values actually mean in terms of extrapolation and can help choose them according to their domain knowledge.
@@
The diagnostic for the CDRC using this kernel is plotted in Figure 6(a). For each mid-dose concentration in both week 6 and 36, the boxplots show how much data support there is in order to calculate the effect of said concentrations on the sample. As the lower limit of the currently used target concentration window is 1 mg/L, the CDRC around that value is especially interesting. We investigate all observations which had a low diagnostic score at concentration 1, as we expect a high bias for these predictions due to data sparsity. Only four observations have a score below 4 effective data points, all of them with a score $<1.4$ EDP. Out of these four observations, three are non-adherent. As these are also the only observations who are classed as non-adherent in the complete cases sample, this is the main reason why the diagnostic score is so low for these individuals. This first sparsity issue identified cannot easily be combated by changing the treatment scheme, as the sparsity is induced not by the treatment variable, but a variable of the adjustment set. And since there are only three data points which share this value, not much can be done to improve the situation. A solution would be to collect more data, which is costly and not usually feasible, or to restrict $L$ by removing adherence from the adjustment set, resulting in a trade-off between bias due to sparsity and bias due to confounding. However, for this last option the magnitude of the induced bias is unclear. Another possible solution is to change the estimand by excluding non-adherent individuals. This leads to the changed diagnostic plot shown in Figure 6(b). The three lowest scoring observations in this new analysis are the three individuals with the highest weight, so if the data support shown by the diagnostic is still deemed unsatisfactory, individuals of higher weight could be excluded from the estimand in a further step. While the diagnostics show that both of these changes in estimand improve data support, the clinical relevance of these estimands has to be judged by medical experts. In general, the diagnostic can help develop strategies to combat sparsity issues based on the data, but whether these strategies are actionable has to be evaluated by domain experts in a last step.
@@ 5.2.2 Causal Analysis with effect modifiers
This plot shows that slow metabolizers are the root of the sparsity problem. As their metabolism works more slowly than for other individuals, they tend to have rather high mid-dose concentrations and there is little support for them at the 1 mg/L mark. This is a typical example of a positivity violation in the treatment variable. Assuming that a slow metabolizer received an efavirenz treatment in accordance with the trial guidelines, which are in turn based on the WHO guidelines, it is not possible for their metabolism to reach a concentration that low after just 12 hours. Therefore, researchers asking the question of “what if” would make a mistake, as that is unrealistic to happen. This problem can be combated by switching the treatment scheme, for example to a dynamic scheme, where slow metabolizers are set to a concentration of 4 mg/L and extensive and intermediate metabolizers to a concentration of 1 mg/L, or a modified treatment scheme, where, again, intermediate and extensive metabolizers receive a concentration of 1 mg/L, but slow metabolizers just receive their natural treatment. Alternatively, the population of interest can be changed by excluding slow metabolizers from the study. Medical experts again have to determine, which of these options would be clinically relevant. For more information on the clinical background see Schomaker et al. [18].
@@ 5.3 Outlook: Multiple Time Point Analysis
The longitudinal case goes beyond the scope of this paper, but we want to present a short overview on how longitudinal data could be approached. For this analysis we used the CHAPAS-3 trial data again. For the longitudinal case we illustrate the estimator-focused diagnostic. As there is no doubly robust estimator for a continuous treatment over multiple time points so far, only IPTW and the g-formula can be used. As IPTW is known to handle extrapolation badly and the sample size is small, we focus on the g-formula here. For the diagnostic hyperparameters, we keep using a Gaussian kernel and the half distance to one standard deviation for the adjustment set and half a standard deviation for the treatment variables for the estimation of the outcome. In this example, we are interested in the estimation of two static intervention schemes. Firstly, $\bar{A}=(1,1,1,1)$ and secondly $\bar{A}=(1,1,10,1)$, where $\bar{A}$ describes the exposure at each of the four time points. The goal is to get an indication of how spikes in the concentration of efavirenz might change the outcome.
@@ 6 Discussion
We developed a diagnostic for positivity issues and other types of sparsity and demonstrated its effectiveness in a simulation study as well as a practical data analysis. We built the diagnostic such that it is first fixed how far extrapolation is possible and then the number of (effective) points for each observation is used to make an informed decision on the trade-off between sparsity induced bias and closeness to the estimand of interest. In theory, the diagnostic can also be constructed the other way round, i. e., one first decides on how many data points are necessary for satisfactory estimation and then the diagnostic shows how far one has to extrapolate for each point in order to achieve that. However, we feel the approach we chose is more intuitive.
@@
There are several possible extensions to this diagnostic measure. The choice of the kernel impacts the results, but it is unclear how to choose a kernel which describes well how all estimators included (for example in the Superlearner) extrapolate. We used a Gaussian kernel in this work, but the viability of other kernel should be explored. In addition, the added value of a data point is not yet considered. For example, it is intuitively clear that several data points which are similar in covariates provide less information on $o^{int}$ compared to the same amount of data points distributed in all directions around $o^{int}$. Therefore, the diagnostic could assume there is a lot of information at a certain point even though there is just a massive amount of information in one particular area close to it and no other data to guide in which way the estimated surface should bend in the direction of the point of interest. There are some possible workarounds. For example one could use a Kullback-Leibler divergence between a density estimate of the weighted points and the kernel in order to gauge whether the points are well distributed within the kernel. However this just adds to the complexity of the diagnostic and is still crude and not founded in any theory which would lead to an actual added value. More research is necessary here. Another possible extension concerns high-dimensional data. The diagnostic might not be useful the way it is presented here, as no data points similar in all covariates can be found for a given observations. For example, there might be a very close patient sharing all 100 variables considered, except gender. The diagnostic would assume there is nothing to be learned from this patient. In order to combat this, the diagnostic can use a harmonic mean of the single dimension diagnostic values or set minimum values for every dimension of the data, so that a difference in gender will at most discount the diagnostic by, for example, 20 %.
One might wonder why the diagnostic is not based on a density estimation of $g_{n}$. Many methods already exist, which use the density in order to decide on whether there is a positivity violation or not. However, we find this only makes sense in a discrete setting and, in addition, a question about sparsity is necessarily a question about data, not the data generating process behind it or the potential of data. Although density estimation is highly difficult, one could build the diagnostic on a type of density estimation instead of pure data, but one would still have to deliberately design this to take the possibility of extrapolation into account. In addition, the estimate of the density has an uncertainty that the data do not have and is less interpretable than the diagnostic we provide.
@@ 7 Conclusion
Sparsity, especially positivity violations, are a big problem for causal effect estimation, but it is rarely discussed how to diagnose the problem. A commonly presented solution is to move the treatment scheme to a (longitudinal) modified treatment policy. However, MTPs are an extremely large class of estimands and it is not clear which MTPs will suffer from sparsity issues and to what extent. In fact, as our simulation study highlighted switching to an MTP is not a panacea and depending on the exact treatment scheme employed, different positivity assumptions need to be fulfilled. The diagnostic we presented provides a simple way of diagnosing sparsity issues for a given estimand. The question of sparsity is a $p$-dimensional problem and the diagnostic reduces this to a one-dimensional output, while also giving the researcher enough information to make the trade-off between estimation bias and closeness to the original research question. In addition, the diagnostic is relatively simple and provides a lot of insight into estimability without having to estimate a single model.
@@ A.1 Extending The Positivity Assumption To Non-Static Interventions
“if $f_{L}(l)\neq 0,f_{A|L}(a|l)>0$ for all a and l”
[10, p. 255]. However, it is unclear what “all $a$ and $l$” are in this context, which is why we want to refine this definition here. “All $l$” refers to the adjustment set of the population considered in the research question. If the population only includes women, the positivity assumption does not need to hold for men. In cases where $l$ has more than one dimension, each dimension can be considered separately. For example, if age and height are included in $l$, there is no need to consider that babies will not be as tall as adults. The condition on the left hand side $f_{L}(l)\neq 0$ is not fulfilled for 190 cm tall newborns, therefore the right hand side does not need to be fulfilled either. The values for “ all $a$” depend on the research question. In order to estimate the ATE for a given value of $a$, the positivity assumption only needs to be fulfilled for this one $a$. In order to estimate a causal dose response curve in the interval $\left[c,d\right]$, the assumption needs to be fulfilled for the whole interval.
@@ A.1.1 This Equals MTP Positivity
$\text{if }(a_{t},h_{t})\in\operatorname{supp}\{A_{t},H_{t}\}\text{ then }(\mathbbm{d}(a_{t},h_{t}),h_{t})\in\operatorname{supp}\{A_{t},H_{t}\}\text{ for}t\in{1,...,\tau}$[8, p. 849]
@@
This shows the only difference in the formulation of the positivity assumption for a static treatment and an MTP is that $a^{\text{obs}}$ is included on the left hand side, as in MTP schemes the intervention can depend on this observed value. The additional $a^{\text{obs}}$ is a relaxation of the positivity assumption (like any additional $l$ would relax it as well) as this can only reduce the the density on the left hand side: $f_{A,L}(a^{\text{obs}},l)=f_{L}(l)\cdot f_{A|L}(a^{\text{obs}}|l)\leq f_{L}(l).$ Compared to a static scheme, the MTP also has the advantage that not every observation needs to receive the same intervention, which also leaves room for the researcher to circumvent positivity issues. These two reasons are why switching to an MTP can help if the originally considered intervention scheme cannot be estimated due to positivity problems.
Appendix B Additional Simulations
B.1 Polynomial

### v3

@@ 5.2 Single Time Point Analysis
At first, in order to simplify the problem at hand, the question is reduced to a single time point target. The variable set is therefore reduced to the initial weight, general adherence, and, if effect modification is of interest, the metabolization group. In addition, the mid-dose concentrations at weeks 6 and 36 are the exposure of interest, which we set to be the same at both time points in the intervention scheme for simplicity. The target is the viral failure measured at week 36 in the form of a dose-response curve. While $a$ can take higher values, we restrict the intervention to a medically more relevant window of $a^{int}\in[0,6]$ [21]. The estimand is therefore $P(V\!F_{t=36}^{EFV_{t=6}=EFV_{t=36}=a^{int}}=1)$ for $a^{int}\in[0,6]$, so the dose-response curve of viral failure in week 36 had the mid-dose concentration of efavirenz been $a^{int}$ in both weeks 6 and 36. In the following, the diagnostic is used in order to analyze how much stochastic positivity/sparsity issues influence the estimation of a CDRC. This is done first for a general causal estimate and then in a second analysis for a causal estimate which includes the estimation of effect modification by the metabolization group.
@@ 5.2.1 Causal Analysis
Before the diagnostic can be used, the hyperparameters have to be fixed. We choose a Gaussian kernel like in the simulation study above. Next, the hyperparameters controlling how wide
@@
Figure 5: The kernel density with its chosen hyperparameters for an $o^{int}$, with weight 3 and mid-dose concentration 5 plotted against the observed data points
@@
or narrow the kernel is in every direction need to be set. Specifically, the question “How far do we think we can extra- and interpolate in each dimension?” has to be asked. For the binary variable adherence, we assume that no extrapolation is possible. For the other adjustment variable weight, we assume that an observation which shares every other variable, but deviates in weight by one standard deviation, should get a diagnostic value of 0.5, meaning it would be treated as half an observation of support. As mid-dose concentration is the exposure variable, we set a more narrow shape in that direction with only half a standard deviation leading to a half observation of support. We can illustrate how that kernel will influence the diagnostic by assuming we are interested in ascribing a concentration of 5 to a person with weight score 3, see Figure 5. If we only consider these two dimensions, this leads to a diagnostic value of roughly 4.3, which is mostly due to the 5 points close by. Plots like this can also help researchers gain a deeper understanding on what certain hyperparameter values mean in terms of extrapolation, and can help choose them according to their domain knowledge.
@@
The diagnostic for the CDRC using this kernel is plotted in Figure 6(a). For each mid-dose concentration in both week 6 and 36, the boxplots show how much data support there is in order to calculate the effect of said concentrations on the sample. As the lower limit of the currently used target concentration window is 1 mg/L, the CDRC around that value is especially interesting. We investigate all observations which had a low diagnostic score at concentration 1, as we expect a high bias for these predictions due to data sparsity. While it is difficult to decide what value of EDP is considered too low, only four observations have a score below 4 effective data points, all of them with a score <1.4<1.4 EDP. Therefore, these four points present clearly as outliers in the diagnostic. Considering the sample size was only 58, expecting an EDP>4>4 for all observations seems not feasible, while an EDP of around 1 is clearly not ideal. These considerations, while still subjective, made it rather clear which observations should be seen as potentially problematic for estimation and are therefore further investigated. Out of these four observations, three are non-adherent. As these are also the only observations who are classed as non-adherent in the complete cases sample, this is the main reason why the diagnostic score is so low for these individuals. This first sparsity issue identified cannot easily be combated by changing the treatment scheme, as the sparsity is induced not by the exposure variable, but a variable of the adjustment set. And since there are only three data points which share this value, not much can be done to improve the situation. A solution would be to collect more data, which is costly and not usually feasible, or to restrict $L$ by removing adherence from the adjustment set, resulting in a trade-off between bias due to sparsity and bias due to confounding. However, for this last option the magnitude of the induced bias is unclear. Another possible solution is to change the estimand by excluding non-adherent individuals. This leads to the changed diagnostic plot shown in Figure 6(b). The three lowest scoring observations in this new analysis are the three individuals with the highest weight, so if the data support shown by the diagnostic is still deemed unsatisfactory, individuals of higher weight could be excluded from the estimand in a further step by setting the target population to adherent observations with a weight below a certain threshold. While the diagnostics show that both of these changes in estimand improve data support, the clinical relevance of these estimands has to be judged by medical experts. In general, the diagnostic can help develop strategies to combat sparsity issues based on the data, but whether these strategies are actionable has to be evaluated by domain experts in a last step.
@@ 5.2.2 Causal Analysis with effect modifiers
This plot shows that slow metabolizers are the root of the sparsity problem. As their metabolism works more slowly than for other individuals, they tend to have rather high mid-dose concentrations and there is little support for them at the 1 mg/L mark. This is a typical example of a positivity violation in the exposure variable. Assuming that a slow metabolizer received an efavirenz treatment in accordance with the trial guidelines, which are in turn based on the WHO guidelines, it is not possible for their metabolism to reach a concentration that low after just 12 hours. Therefore, researchers asking the question of “what if” would make a mistake, as that is unrealistic to happen. This problem can be combated by switching the exposure scheme, for example to a dynamic scheme, where slow metabolizers are set to a concentration of 4 mg/L and extensive and intermediate metabolizers to a concentration of 1 mg/L, or a modified treatment scheme, where, again, intermediate and extensive metabolizers receive a concentration of 1 mg/L, but slow metabolizers just receive their natural exposure. Alternatively, the population of interest can be changed by excluding slow metabolizers from the study. Medical experts again have to determine, which of these options would be clinically relevant. For more information on the clinical background see Schomaker et al. [21].
@@ 5.3 Outlook: Multiple Time Point Analysis
The longitudinal case goes beyond the scope of this paper, but we want to present a short overview on how longitudinal data could be approached. For this analysis we used the CHAPAS-3 trial data again. For the longitudinal case we illustrate the estimator-focused diagnostic. The goal here is to estimate the CDRC for a continuous treatment measured over multiple time points. While doubly robust estimands, for which doubly robust estimators can be constructed, exist for this data scenario, like the LMTP [9], it is not possible to construct one for the CDRC [26, 27]. Therefore, only IPTW and the g-formula can be used here. As IPTW is known to handle extrapolation badly and the sample size is small, we focus on the g-formula here. For the diagnostic hyperparameters, we keep using a Gaussian kernel and the half distance to one standard deviation for the adjustment set and half a standard deviation for the exposure variables for the estimation of the outcome. In this example, we are interested in the estimation of two static intervention schemes. Firstly, $\bar{A}=(1,1,1,1)$ and secondly $\bar{A}=(1,1,10,1)$, where $\bar{A}$ describes the exposure at each of the four time points. The goal is to get an indication of how spikes in the concentration of efavirenz might change the outcome.
@@ 6 Discussion
We developed a diagnostic for stochastic positivity issues and other types of sparsity and demonstrated its effectiveness in a simulation study as well as a practical data analysis. We built the diagnostic such that it is first fixed how far extrapolation is possible and then the number of (effective) points for each observation is used to make an informed decision on the trade-off between sparsity induced bias and closeness to the estimand of interest. In theory, the diagnostic can also be constructed the other way round, i. e., one first decides on how many data points are necessary for satisfactory estimation and then the diagnostic shows how far one has to extrapolate for each point in order to achieve that. However, we feel the approach we chose is more intuitive.
@@
There are several possible extensions to this diagnostic measure. The choice of the kernel impacts the results, but it is unclear how to choose a kernel which describes well how all estimators included (for example in the Superlearner) extrapolate. We used a Gaussian kernel in this work, but the viability of other kernels should be explored. In addition, the added value of a data point is not yet considered. For example, it is intuitively clear that several data points which are similar in covariates provide less information on $o^{int}$ compared to the same amount of data points distributed in all directions around $o^{int}$. Therefore, the diagnostic could assume there is a lot of information at a certain point even though there is just a massive amount of information in one particular area close to it and no other data to guide in which way the estimated surface should bend in the direction of the point of interest. There are some possible workarounds. For example one could use a Kullback-Leibler divergence between a density estimate of the weighted points and the kernel in order to gauge whether the points are well distributed within the kernel. However, this just adds to the complexity of the diagnostic and is still crude and not founded in any theory which would lead to an actual added value. More research is necessary here. Another possible extension concerns high-dimensional data. In such a case, the diagnostic might not be useful the way it is presented here, as no data points similar in all covariates can be found for a given observation. For example, there might be a very close patient sharing all 100 variables considered, except gender. The diagnostic would assume there is nothing to be learned from this patient. In order to combat this, the diagnostic can use a harmonic mean of the single dimension diagnostic values or set minimum values for every dimension of the data, so that a difference in gender will at most discount the diagnostic by, for example, 20 % (for more see Appendix B).
One might wonder why the diagnostic is not based on a density estimation of $g_{n}$. Many methods already exist that use the density in order to decide on whether there is a stochastic positivity violation or not. However, we find this only makes sense in a discrete setting and, in addition, a question about sparsity is necessarily a question about data, not the data generating process behind it or the potential of data. Density estimation is highly difficult, and although one could build the diagnostic on a type of density estimation instead of pure data, but one would still have to deliberately design this to take the possibility of extrapolation into account. In addition, the estimate of the density has an uncertainty that the pure data does not have and we feel it would be less interpretable than the diagnostic we provide.
@@ 7 Conclusion
Sparsity, especially in the form of stochastic positivity violations, is a big problem for causal effect estimation, but it is rarely discussed how to identify the problem. A commonly presented solution is to move the treatment scheme to a (longitudinal) modified treatment policy. However, MTPs are an extremely large class of estimands and it is not clear which MTPs will suffer from sparsity issues and to what extent. In fact, as our simulation study highlighted, switching to an MTP is not a panacea and depending on the exact treatment scheme employed, different positivity assumptions need to be fulfilled. The diagnostic we presented provides a simple way of diagnosing sparsity issues for a given estimand. It works for every kind of estimand, including static schemes, like a Causal Dose Response Curve, dynamic schemes, and Modified Treatment Policies, by comparing the observed data and the dataset under intervention. The question of sparsity is a $p$-dimensional problem and the diagnostic reduces this to a one-dimensional output, while also giving the researcher enough information to make the trade-off between estimation bias and closeness to the original research question. In addition, the diagnostic is relatively simple and provides a lot of insight into estimability without having to estimate a single model.
@@ A.1 Extending The Positivity Assumption To Non-Static Interventions
“if $f_{L}(l)\neq 0,f_{A|L}(a|l)>0$ for all $a$ and $l$
[13, p. 255]. However, it is unclear what “all $a$ and ll” are in this context, which is why we want to refine this definition here. “All ll” refers to the adjustment set of the population considered in the research question. If the population only includes women, the positivity assumption does not need to hold for men. In cases where $l$ has more than one dimension, each dimension can be considered separately. For example, if age and height are included in $l$, there is no need to consider that babies will not be as tall as adults. The condition on the left hand side $f_{L}(l)\neq 0$ is not fulfilled for 190 cm tall newborns, therefore the right hand side does not need to be fulfilled either. The values for “ all aa” depend on the research question. In order to estimate the ATE for a given value of $a$, the positivity assumption only needs to be fulfilled for this one $a$. In order to estimate a causal dose response curve in the interval $\left[c,d\right]$, the assumption needs to be fulfilled for the whole interval.
@@ A.1.1 This Equals MTP Positivity
$\text{if }(a_{t},h_{t})\in\operatorname{supp}\{A_{t},H_{t}\}\text{ then }(\mathbbm{d}(a_{t},h_{t}),h_{t})\in\operatorname{supp}\{A_{t},H_{t}\}\text{ for }t\in{1,...,\tau}$[9, p. 849]
@@
This shows the only difference in the formulation of the positivity assumption for a static treatment and an MTP is that $a^{\text{obs}}$ is included on the left hand side, as in MTP schemes the intervention can depend on this observed value. The additional $a^{\text{obs}}$ is a relaxation of the positivity assumption (like any additional $l$ would relax it as well) as this can only reduce the density on the left hand side: $f_{A,L}(a^{\text{obs}},l)=f_{L}(l)\cdot f_{A|L}(a^{\text{obs}}|l)\leq f_{L}(l).$ Compared to a static scheme, the MTP also has the advantage that not every observation needs to receive the same intervention, which also leaves room for the researcher to circumvent positivity issues. These two reasons are why switching to an MTP can help if the originally considered intervention scheme cannot be estimated due to positivity problems.
Appendix B The Curse Of Dimensionality
Figure 8: Simulation data showing the change in EDPEDP for a higher dimensionality of the adjustment set.
High-dimensional data present a difficulty in estimation which is reflected in the Effective Data Points. In order to illustrate the behavior of the diagnostic for higher dimensional adjustment sets, we present an additional simulation here. The data is simulated as follows:
$A$ $\sim\mathrm{Normal}(0,1)$
$L_{p}$ $\sim\mathrm{Normal}(0,1),$
with p=0,,9p=0,...,9 and $P$, the number of variables included in the adjustment set, incremented for every calculation of the diagnostic. The intervention strategy used for the diagnostic was a simple static treatment of $A=0$ for all 1000 observations. Figure 8 shows that the EDP drops drastically with a higher dimensionality of the adjustment set. This represents the Curse Of Dimensionality: the more data dimensions there are the more the information is “spread out” meaning with more features data points are bound to get more dissimilar from each other. However, this simulation study represents a sort of worst case scenario, where there is no dependence between the dimensions. In realistic scenarios data are expected to be more clustered, e. g. in human body measurement data one would expect people with longer arms to have longer legs also. Therefore, the real drop in EDP will likely be less dramatic than in this example, but in general higher dimensions lead to lower EDP.
When working with high-dimensional data, this problem is usually dealt with by making additional assumptions on how complex the functional shape of the predictive model can be, for example by limiting the number of interaction effects. If these assumptions are clearly necessary due to the dimensionality of the data compared to the sample size, they should be reflected in the diagnostic. Often, however, these assumptions are left to the algorithms. In order to still obtain meaningful insights from the data, we propose two adjustments to the diagnostic. In both cases, the kernel function is applied separately to each dimension (so the one-dimensional Gaussian if a multivariate Gaussian is utilized), and then the results are combined in a modified way. The minimum variant prevents any single variable from too drastically reducing the overall diagnostic. Specifically, we cap the diagnostic value for each dimension $EDP_{p}$ at a certain minimum, and then compute the overall EDP as
$\text{EDP}=\prod_{p=1}^{P}\max(\text{EDP}_{p},\text{minval}_{p})$
Here, $minval_{p}$ is a minimum value that can be chosen based on how important variable $p$ is expected to be. For example, for the most important confounders no minimum value is set, while for a confounder which is expected to only have a minor influence on treatment and outcome, a minimum value of $0.5$ could be set. If it’s unclear how to set these values, because there is little information on the importance of each feature, the second method may be preferable. Instead of using a product, the harmonic mean variant combines the per-dimension diagnostics using the harmonic mean:
$\text{EDP}=\left(\frac{1}{P}\sum_{p=1}^{P}\frac{1}{\text{EDP}_{p}}\right)^{-1}$
This has the effect of reducing the impact of very low values without requiring subjective minimum thresholds.
Appendix C Additional True Functional Shapes For Simulation
C.1 Polynomial